# Modelling reliability of reversible circuits with 2D second-order cellular automata

### 20 Dec 2023 v1

@@ 3.1 Damage propagation from single defect
Quantification of damage propagation in CA can be useful for reliability analysis of logic circuits based on such a model. Informal idea of Lyapunov exponents sometimes used as a measure of such propagation [20] and few approaches to rigorous definition are developed [34, 35, 36, 37, 38, 39] despite of certain difficulties and ambiguities related with application of such quantity to discrete systems. An intensity of defect accumulation and speed of the damage spread could be mentioned as two alternative views on the Lyapunov exponent [38].
Let us note some subtleties with definition of exponential quantities in CA. For a single defect even for an extreme case of influence on all neighbors after first step for $d$-dimensional CA there are $\epsilon=3^{d}$ affected cells in Moore neighborhood, but number of such cells grows only polynomially with time $\epsilon_{t}=(3t)^{d}$ and this amount is the maximally possible for any CA with local update rule. It provides natural constraint on exponential growth if to count defects as the total number of affected cells
$\epsilon_{t}=O(e^{\lambda t}),\quad\lambda=\lim_{t\to\infty}\frac{1}{t}\ln\epsilon_{t}\leq\lim_{t\to\infty}\frac{1}{t}\ln(3t)^{d}=\lim_{t\to\infty}\frac{1}{t}(\ln 3+\ln t)d=0\,.$ (8)
Such a problem could be addressed by consideration of separate replica of a system for each defect for any time step [34] with formal use of so-called Boolean derivatives [40], but such calculation without proper accounting of defects cancellation may produce overestimation of possible effects.
@@
where $\oplus$ is addition modulo 2. According to enumeration in Ref. [41] the CA corresponds to rule 90 and provides important example of additive CA [42]. Statistical properties of second-order RCA produced by such a rule denoted as 90R was discussed in Ref. [43].
For initial (irreversible) rule 90 CA calculation with Boolean derivatives produces value $\lambda=\ln 2$ for Lyapunov exponent [34]. For configuration with single occupied cell the total number of non-empty cells after $n$ steps is described by sequence A001316 in OEIS [44] and can be expressed as $2^{wt({n})}$, where $wt({n})$ is number of units in binary expansion of $n$ (Hamming weight).
Despite of exponential notation the value $2^{wt({n})}$ may not exceed $n+1$, but for any even $n$ the value doubles on next step $n+1$ resembling dynamics that could be expected for process with $\lambda=\ln 2$. Damage front for such CA moves with unit speed in both directions and could provide only minimal information for comparison with other CA [34].
@@
This CA is notable as a simple example with self-replication of any pattern discovered by E. Fredkin about 1960 and recollected by M. Gardner [25, Ch. 21]. For a single occupied cell after $n$ steps there are $4^{wt({n})}$ non-empty cells, see sequence A102376 in OEIS. For particular values of $t$ the configurations correspond to the square grids with different number of points and intervals between them. Due to Eq. (10) the CA is linear with respect to addition modulo 2 and because any pattern can be formally represented as a sum of few configurations with single occupied cell, at appropriate time steps the pattern is replicated over grids with sufficiently large intervals.
Using earlier conventions for CEOT rules such CA could be formally denoted as $C_{0-4}E_{1,3}$, but the notation $E_{1,3}$ is more natural for this CA with von Neumann neighborhood, because the local rule Eq. (10) does not depend on number of cells with common corner. The CA should be distinguished from similar replicator with Moore neighborhood [45], see sequence A160239 in OEIS.
Second-order 2D RCA from $E_{1,3}$ can be denoted as $E_{1,3}^{I\!I}$. Number of cells with different states after $n$ steps for configuration with a single cell with state 1 (red) is described by sequences A244642 and A244643 in OEIS. The RCA is not replicator, but it may produce interesting fractal-like patterns. For configuration with single red cell evolution is also coincides with considered earlier RCA $C_{0}E_{1}^{\,I\!I}$ [31] and for $C_{0}E_{1,3}^{\,I\!I}$ such coincidence also can be shown in a similar way due to some similarity of local rules with $C_{0}E_{1}^{\,I\!I}$ and $E_{1,3}^{I\!I}$, see Fig. 6.
@@
The numbers $N_{1}$ and $N_{2}$ of cells with states 1 and 2 after $n$ steps for configuration with single red cell satisfy recursive formulas [31]
@@
The first part of each equation would correspond to asymptotically quadratic growth $O(n^{2})$, but the second part to only linear one $O(n)$. Such a variance in formal estimation of dimensions seems to be in some agreement with the shape of configurations Fig. 6 resembling some geometrical sequences used for construction of fractals [46].
Such behavior justifies consideration of models with polynomial growth together with an analogy of formal or fractal dimension [46] already investigated earlier in context of CA [18, Ch. 4] and denoted further as ${\mathfrak{D}}$ that is not vanishing unlike $\lambda$ Eq. (8)
@@
The subtleties of transition from definition of fractal dimension relevant to continuous case to such discrete systems as CA can be found in Ref [18, Ch. 4], but here formal dimension ${\mathfrak{D}}$ is treated rather as convenient quantity defined by Eq. (12).
@@
The evolution of patterns from a single red cell for other RCA can be quite simple with ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ and $C_{0}E_{1-3}^{\,I\!I}$ depicted on Fig. 7 and ${\mathfrak{D}}_{\infty}=1$ for $C_{0}E_{+}^{\,I\!I}$ depicted on Fig. 8.
@@
Evolution of isotropic RCA $C_{0}E_{1,2o}^{\,I\!I}$ from single red cell is shown on Fig. 9. Analytical expression for number of cells is not known to author, but numerical estimation of ${\mathfrak{D}}^{\diamond}_{t}$ shows slow growth with minimal variations and upper limit by some value ${\mathfrak{D}}^{\diamond}_{\max}\leq 2$. For example, ${\mathfrak{D}}^{\diamond}_{\max}$ for $C_{0}E_{1,2o}^{\,I\!I}$ is increasing from ${\mathfrak{D}}^{\diamond}\simeq 1.55$ for $t=2$ to ${\mathfrak{D}}^{\diamond}\simeq 1.88$ for $t=1000$.
@@
However, behavior becomes less trivial for more complex configurations. For example static $2\times 2$ blue block added to initial configuration with single red cell does not change values ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ and ${\mathfrak{D}}_{\infty}=1$ for $C_{0}E_{+}^{\,I\!I}$, but for $C_{0}E_{1-3}^{\,I\!I}$ growing rectangular area of empty cells appears inside of initial diamond patters resulting ${\mathfrak{D}}_{\infty}=1$.
Behavior looks less predictable if to add static configurations with more complex shapes. In such a case ${\mathfrak{D}}^{\diamond}$ for $C_{0}E_{+}^{\,I\!I}$ can increase from ${\mathfrak{D}}_{\infty}=1$ to ${\mathfrak{D}}_{\infty}=2$ due to many interfering damage fronts and for $C_{0}E_{1-3}^{\,I\!I}$ earlier mentioned decrease from ${\mathfrak{D}}_{\infty}=2$ to ${\mathfrak{D}}_{\infty}=1$ also can be suppressed due to similar effects. In both cases limiting value may depend in some non-trivial way from a shape of static configuration and initial position of red cell.
@@ 3.2 Damage from improper signals interaction
For example, all signals should enter in a gate at the same time. A few kinds of errors could be expected due to delay between signals. The delay of output signals could be considered as a minimal possible problem. The simple example is the swap gate Fig. 1. However, even for swap gates there is critical range of delays producing more serious problems. For ‘wide’ swap gate such a critical range for delay of second (lower) signal is $3-11$ time steps and for ‘narrow’ swap the range is even bigger ($3-15$), because signal trajectories intersect repeatedly.
@@
Collisions of signals with different delays are shown for discussed in this work RCA on Figs. 1014. Due to symmetry of picture with respect to reflections about the diagonal only six values $0\leq|\Delta|\leq 5$ are shown for each case. The value $|\Delta|=5$ is the minimal one without any interaction.
Figure 10: Collisions with different delays in $C_{0}E_{1}^{\,I\!I}$
Figure 11: Collisions with different delays in $C_{0}E_{1,3}^{\,I\!I}$
Figure 12: Collisions with different delays in $C_{0}E_{1,2}^{\,I\!I}$
Figure 14: Collisions with different delays in $C_{0}E_{1,2o}^{\,I\!I}$
For RCA $C_{0}E_{1}^{\,I\!I}$ and $C_{0}E_{1,3}^{\,I\!I}$ of the first type sometime referred earlier as ‘snakes’ depicted on Fig. 10 and Fig. 11 the value $\Delta=0$ corresponds to reflection of signals followed by motion in opposite direction. The values $1\leq|\Delta|\leq 3$ produce a signal expanding in two opposite directions together with initial one. The value $|\Delta|=4$ seems produce more critical damage with front expanding in all possible directions.
The same value $|\Delta|=4$ also looks critical for second type of RCA $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ sometimes referred earlier as ‘ants’ depicted on Fig. 12 – Fig. 14, but only for $C_{0}E_{1,2o}^{\,I\!I}$ value $\Delta=0$ produces simple reflection and for other four RCA of such type $\Delta=0$ causes more critical omnidirectional damage spread. The values $1\leq|\Delta|\leq 3$ for all four RCA of such type due to split of one signal after collision produce three signals moving apart with apparently smaller damage.
Both $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ produce the same damage patterns Fig. 13 with ${\mathfrak{D}}_{\infty}=2$ for $\Delta=0$ or ${\mathfrak{D}}_{\infty}=1$ for $|\Delta|=4$, despite of different behavior for initial configuration with single red cell discussed earlier, see Fig. 7 and Fig. 8.
The properties of simple collisions may only provide some guesses about behavior in more complex elements and computer simulations are necessary to clarify details. For already mentioned ‘wide’ swap the delay 7 time steps between two signals corresponds to $\Delta=0$ and produces reflection with reverse motion of signals for $C_{0}E_{1}^{\,I\!I}$, $C_{0}E_{1,3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$, but for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ the damage is more serious resembling behavior for boundary values 3 and 11 of critical range.
@@
Intersection of signals in swap gates or other kinds of routing can be considered as particular problem of 2D design and resolved by analogue 3D RCA models. However for any non-trivial gate discussed above interaction of signals is necessary process. Let us consider elementary useful interaction relevant already for simple switch gate Fig. 2 to investigate errors due to wrong delays.
Such interactions are depicted on Figs. 1519. The main difference in comparison with Figs. 1014 is $2\times 2$ static blue block (cells with state 3) for proper signal motion after collision. The interaction between signals again corresponds to $-4\leq\Delta\leq 4$ and six values $0\leq|\Delta|\leq 5$ are enough for each case due to symmetry with respect to reflections about the diagonal.
Figure 15: Interactions with different delays in $C_{0}E_{1}^{\,I\!I}$
Figure 16: Interactions with different delays in $C_{0}E_{1,3}^{\,I\!I}$
Figure 17: Interactions with different delays in $C_{0}E_{1,2}^{\,I\!I}$
Figure 19: Interactions with different delays in $C_{0}E_{1,2o}^{\,I\!I}$
For $1\leq|\Delta|\leq 3$ after interaction two signals move along parallel paths with relative delay $|\Delta|-1$. For elements similar with switch gate Fig. 2 proper value is $|\Delta|=1$ taking into account necessity for repeated interaction. The value $|\Delta|=4$ produces omnidirectional damage front for all RCA discussed in the paper and $\Delta=0$ is related with similar problem for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$. For $C_{0}E_{1}^{\,I\!I}$, $C_{0}E_{1,3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ the same value $|\Delta|=0$ causes reflection of both signals after interaction.
Both $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ produce the same damage fronts with ${\mathfrak{D}}_{\infty}=1$, see Fig. 18. Despite of some resemblance for $\Delta=0$ of initial time steps shown on Fig. 17, ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ not only for $|\Delta|=4$, but also for $\Delta=0$ with shape of damage described as diamond-like area with rectangular empty space due to initial impact of $2\times 2$ static blue block.
@@ 3.3 Consequences of damage distribution
For ${\mathfrak{D}}_{\infty}=1$ omnidirectional damage front (see Fig. 8) also may produce the similar disruption of functionality. An alternative example of ${\mathfrak{D}}_{\infty}=1$ with anisotropic distribution can be illustrated by result of collision RCA $C_{0}E_{1}^{\,I\!I}$ or $C_{0}E_{1,3}^{\,I\!I}$ (‘snakes’) with static block. In such a case the signal starts expansion in two opposite direction. If only one of two ‘heads’ of expanding line reaches some obstacle further evolution correspond to motion of a signal in unconstrained direction.
@@
However, even for $C_{0}E_{1}^{\,I\!I}$ or $C_{0}E_{1,3}^{\,I\!I}$ already interaction of two signals may produce omnidirectional spreading of damage with ${\mathfrak{D}}_{\infty}=1$ or ${\mathfrak{D}}_{\infty}=2$ (see Fig. 10, Fig. 11, Fig. 15 and Fig. 16 for $|\Delta|=4$) resulting disruption of circuit functionality.
For other RCA such as $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ (‘ants’) even for one signal collision with static block leads to the appearance of two signals moving in opposite directions. Unconstrained motion of such signals formally would result to ${\mathfrak{D}}_{\infty}=0$, but consequent collisions with static elements produce a fast increase in the number of signals. The interaction between such signals again may produce omnidirectional damage spread (see Fig. 12 – Fig. 14 and Fig. 17 – Fig. 19) generating errors in the gates affected by its propagation.
@@
For example, already discussed minimum intervals between consequent signals for ‘wide’ and ‘narrow’ swap gates on Fig. 1 should be 12 and 16 time steps respectively. For Fredkin gate Fig. 3 minimum interval depends on the composition of the signal (representing combinations of zeros and units) in two consequent groups. In the worst case such interval should not be less than 22 time steps. Controlled-NOT gate on Fig. 5 can be used for RCA $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ with a relatively short minimum possible interval of 10 time steps between consequent groups of signals.

### 19 Mar 2024 v2

@@ 3.1 Damage propagation from single defect
Quantification of damage propagation in CA can be useful for reliability analysis of logic circuits based on such a model. Informal idea of Lyapunov exponents sometimes used as a measure of such propagation WolCA and few approaches to rigorous definition are developed BBR92 ; MS92 ; Ti00 ; BB10 ; BG18 ; Ko21 despite of certain difficulties and ambiguities related with application of such quantity to discrete systems. An intensity of defect accumulation and speed of the damage spread could be mentioned as two alternative views on the Lyapunov exponent BG18 .
Let us note some subtleties with definition of exponential quantities in CA. For a single defect even for an extreme case of influence on all neighbors after first step for $d$-dimensional CA there are $\epsilon=3^{d}$ affected cells in Moore neighborhood, but number of affected cells at $t$-th steps denoted further as $\epsilon_{t}$ may grow only polynomially with time $\epsilon_{t}=(2t+1)^{d}$ and this amount is the maximally possible for any CA with local update rule. It provides natural constraint on exponential growth if to count defects as the total number of affected cells
$\epsilon_{t}=O(e^{\lambda t}),\quad\lambda=\lim_{t\to\infty}\frac{\ln\epsilon_{t}}{t}\leq\lim_{t\to\infty}\frac{\ln(2t+1)^{d}}{t}=\lim_{t\to\infty}\frac{d}{t}\ln(2t+1)=0\,.$ (8)
Such a problem could be addressed by consideration of separate replica of a system for each defect for any time step BBR92 with formal use of so-called Boolean derivatives Vi90 , but such calculation without proper accounting of defects cancellation may produce overestimation of possible effects.
@@
where $\oplus$ is addition modulo 2. According to enumeration in Ref. W93 the CA corresponds to rule 90 and provides important example of additive CA AlgCA . Statistical properties of second-order RCA produced by such a rule denoted as 90R was discussed in Ref. Ta87 .
For initial (irreversible) rule 90 CA calculation with Boolean derivatives produces value $\lambda=\ln 2$ for Lyapunov exponent BBR92 . For configuration with single occupied cell the total number of non-empty cells after $n$ steps is described by sequence A001316 in OEIS OEIS and can be expressed as $2^{wt({n})}$, where $wt({n})$ is number of units in binary expansion of $n$ (Hamming weight).
Despite of exponential notation the value $2^{wt({n})}$ may not exceed $n+1$, but for any even $n$ the value doubles on next step $n+1$ resembling dynamics that could be expected for process with $\lambda=\ln 2$. Damage front for such CA moves with unit speed in both directions and could provide only minimal information for comparison with other CA BBR92 .
@@
This CA is notable as a simple example with self-replication of any pattern discovered by E. Fredkin about 1960 and recollected by M. Gardner (Gard, , Ch. 21). For a single occupied cell after $n$ steps there are $4^{wt({n})}$ non-empty cells, see sequence A102376 in OEIS. For particular values of $t$ the configurations correspond to the square grids with different number of points and intervals between them. Due to Eq. (10) the CA is linear with respect to addition modulo 2 and because any pattern can be formally represented as a sum of few configurations with single occupied cell, at appropriate time steps the pattern is replicated over grids with sufficiently large intervals.
Using earlier conventions for CEOT rules such CA could be formally denoted as $C_{0-4}E_{1,3}$, but the notation $E_{1,3}$ is more natural for this CA with von Neumann neighborhood, because the local rule Eq. (10) does not depend on number of cells with common corner. The CA should be distinguished from similar replicator with Moore neighborhood SlCA , see sequence A160239 in OEIS.
Second-order 2D RCA from $E_{1,3}$ can be denoted as $E_{1,3}^{I\!I}$. Number of cells with different states after $n$ steps for configuration with a single cell with state 1 (red) is described by sequences A244642 and A244643 in OEIS. The RCA $E_{1,3}^{I\!I}$ is not replicator, but it may produce interesting fractal-like patterns. For configuration with single red cell evolution is also coincides with considered earlier RCA $C_{0}E_{1}^{\,I\!I}$ AV14 and for $C_{0}E_{1,3}^{\,I\!I}$ such coincidence also can be shown in a similar way due to some similarity of local rules with $C_{0}E_{1}^{\,I\!I}$ and $E_{1,3}^{I\!I}$, see Figure 6.
@@
The numbers $N_{1}$ and $N_{2}$ of cells with states 1 and 2 after $n$ steps for configuration with single red cell satisfy recursive formulas AV14
@@
The first part of each equation would correspond to asymptotically quadratic growth $O(n^{2})$, but the second part to only linear one $O(n)$. Such a variance in formal estimation of dimensions seems to be in some agreement with the shape of configurations Figure 6 resembling some geometrical sequences used for construction of fractals Frac .
Such behaviour justifies consideration of models with polynomial growth together with an analogy of formal or fractal dimension Frac already investigated earlier in context of CA (Ila, , Ch. 4) and denoted further as ${\mathfrak{D}}$ that is not vanishing unlike $\lambda$ Eq. (8)
@@
The subtleties of transition from definition of fractal dimension relevant to continuous case to such discrete systems as CA can be found in Ref (Ila, , Ch. 4), but here formal dimension ${\mathfrak{D}}$ is treated rather as convenient quantity defined by Eq. (12).
@@
The evolution of patterns from a single red cell for other RCA can be quite simple with ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ and $C_{0}E_{1-3}^{\,I\!I}$ depicted on Figure 7 and ${\mathfrak{D}}_{\infty}=1$ for $C_{0}E_{+}^{\,I\!I}$ depicted on Figure 8.
@@
Evolution of isotropic RCA $C_{0}E_{1,2o}^{\,I\!I}$ from single red cell is shown on Figure 9. Analytical expression for number of cells is not known to author, but numerical estimation of ${\mathfrak{D}}^{\diamond}_{t}$ shows slow growth with minimal variations and upper limit by some value ${\mathfrak{D}}^{\diamond}_{\max}\leq 2$. For example, ${\mathfrak{D}}^{\diamond}_{\max}$ for $C_{0}E_{1,2o}^{\,I\!I}$ is increasing from ${\mathfrak{D}}^{\diamond}\simeq 1.55$ for $t=2$ to ${\mathfrak{D}}^{\diamond}\simeq 1.88$ for $t=1000$.
@@
However, behaviour becomes less trivial for more complex configurations. For example static $2\times 2$ blue block added to initial configuration with single red cell does not change values ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ and ${\mathfrak{D}}_{\infty}=1$ for $C_{0}E_{+}^{\,I\!I}$, but for $C_{0}E_{1-3}^{\,I\!I}$ growing rectangular area of empty cells appears inside of initial diamond patters resulting ${\mathfrak{D}}_{\infty}=1$.
Behaviour looks less predictable if to add static configurations with more complex shapes. In such a case ${\mathfrak{D}}^{\diamond}$ for $C_{0}E_{+}^{\,I\!I}$ can increase from ${\mathfrak{D}}_{\infty}=1$ to ${\mathfrak{D}}_{\infty}=2$ due to many interfering damage fronts and for $C_{0}E_{1-3}^{\,I\!I}$ earlier mentioned decrease from ${\mathfrak{D}}_{\infty}=2$ to ${\mathfrak{D}}_{\infty}=1$ also can be suppressed due to similar effects. In both cases limiting value may depend in some non-trivial way from a shape of static configuration and initial position of red cell.
@@ 3.2 Damage from improper signals interaction
For example, all signals should enter in a gate at the same time. A few kinds of errors could be expected due to delay between signals. The delay of output signals could be considered as a minimal possible problem. The simple example is the swap gate Figure 1. However, even for swap gates there is critical range of delays producing more serious problems. For ‘wide’ swap gate such a critical range for delay of second (lower) signal is $3-11$ time steps and for ‘narrow’ swap the range is even bigger ($3-15$), because signal trajectories intersect repeatedly.
@@
Collisions of signals with different delays are shown for discussed in this work RCA on Figures 1 – 5 in A.
The properties of simple collisions may only provide some guesses about behaviour in more complex elements and computer simulations are necessary to clarify details. For already mentioned ‘wide’ swap the delay 7 time steps between two signals corresponds to $\Delta=0$ and produces reflection with reverse motion of signals for $C_{0}E_{1}^{\,I\!I}$, $C_{0}E_{1,3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$, but for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ the damage is more serious resembling behaviour for boundary values 3 and 11 of critical range.
@@
Intersection of signals in swap gates or other kinds of routing can be considered as particular problem of 2D design and resolved by analogue 3D RCA models. However for any non-trivial gate discussed above interaction of signals is necessary process. An elementary interaction relevant already for simple switch gate Figure 2 can be considered to investigate errors due to wrong delays. Such interactions are depicted on Figures 1 – 5 in B.
@@ 3.3 Consequences of damage distribution
For ${\mathfrak{D}}_{\infty}=1$ omnidirectional damage front (see Figure 8) also may produce the similar disruption of functionality. An alternative example of ${\mathfrak{D}}_{\infty}=1$ with anisotropic distribution can be illustrated by result of collision RCA $C_{0}E_{1}^{\,I\!I}$ or $C_{0}E_{1,3}^{\,I\!I}$ (‘snakes’) with static block. In such a case the signal starts expansion in two opposite direction. If only one of two ‘heads’ of expanding line reaches some obstacle further evolution correspond to motion of a signal in unconstrained direction.
@@
However, even for $C_{0}E_{1}^{\,I\!I}$ or $C_{0}E_{1,3}^{\,I\!I}$ already interaction of two signals may produce omnidirectional spreading of damage with ${\mathfrak{D}}_{\infty}=1$ or ${\mathfrak{D}}_{\infty}=2$ resulting disruption of circuit functionality (see Figures 12 in A or Figures 12 in B for $|\Delta|=4$).
For other RCA such as $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ (‘ants’) even for one signal collision with static block leads to the appearance of two signals moving in opposite directions. Unconstrained motion of such signals formally would result to ${\mathfrak{D}}_{\infty}=0$, but consequent collisions with static elements produce a fast increase in the number of signals. The interaction between such signals again may produce omnidirectional damage spread (see Figures 3 – 5 in A and Figures 3 – 5 in B) generating errors in the gates affected by its propagation.
@@
For example, already discussed minimum intervals between consequent signals for ‘wide’ and ‘narrow’ swap gates on Figure 1 should be 12 and 16 time steps respectively. For Fredkin gate Figure 3 minimum interval depends on the composition of the signal (representing combinations of zeros and units) in two consequent groups. In the worst case such interval should not be less than 22 time steps. Controlled-NOT gate on Figure 5 can be used for RCA $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ with a relatively short minimum possible interval of 10 time steps between consequent groups of signals.